# Dynamically Adaptive Reasoning via LLM-Guided MCTS for Efficient and Context-Aware KGQA

### v3

@@ Conclusion
Appendix E E. More experimental results
More Sensitivity Analysis
More case study
Table 7 presents a case study comparing the answer accuracy of DAMR with general LLMs: Llama-2-13B, Qwen-3-14B, GPT 4.1-mini, and GPT 4.1. When asked about the origin of the Greek language, all baseline models generate fluent and seemingly plausible responses grounded in general linguistic knowledge, such as “Proto-Indo-European” or “Proto-Hellenic”, but fail to identify the correct answer: Attic Group. In contrast, DAMR accurately predicts the correct entity by explicitly traversing the relation path base.rosetta.languoid.parent within the knowledge graph. This example illustrates a key advantage of DAMR: rather than relying solely on learned linguistic patterns, it performs structured reasoning over the knowledge graph, enabling precise and faithful answers to ontology-specific queries that often elude general-purpose LLMs.
Appendix F F. Prompt Template
@@
Prompt Template for LLM-Guided Path Expansion Role You are an expert assistant for Knowledge Graph Question Answering (KGQA). Your core capability is to deeply understand natural language questions and the semantics of knowledge graph relations to find the most relevant reasoning paths. Task Your task is to act as a “Relation Retriever.” Given a natural language question and a list of candidate relations, you must analyze the semantics of the question and each relation to select up to k relations that are most likely to lead to the correct answer. Rules and Constraints Fidelity to Candidates: Your selection of relations MUST come strictly from the provided Candidate Relations list. Do not invent or modify relations. Quantity Limit: Return no more than k relations. If multiple relations are highly relevant, order them from most to least relevant. If there are fewer than k relevant relations, return only those. Output Format: Your response MUST be a list of strings, containing the names of the relations you have selected. Example Input: Question: ”who was the president after jfk died” Candidate Relations: {”government.president”, ”government.president.successor”, ”location.location.containedby”, ”people.person.place_of_birth”} K: 2 Output:

### v4

@@ Appendix D More experimental results
Appendix E Prompt Template
@@
Prompt Template for LLM-Guided Path Expansion Role You are an expert assistant for Knowledge Graph Question Answering (KGQA). Your core capability is to deeply understand natural language questions and the semantics of knowledge graph relations to find the most relevant reasoning paths. Task Your task is to act as a “Relation Retriever.” Given a natural language question and a list of candidate relations, you must analyze the semantics of the question and each relation to select up to k relations that are most likely to lead to the correct answer. Rules and Constraints Fidelity to Candidates: Your selection of relations MUST come strictly from the provided Candidate Relations list. Do not invent or modify relations. Quantity Limit: Return no more than k relations. If multiple relations are highly relevant, order them from most to least relevant. If there are fewer than k relevant relations, return only those. Output Format: Your response MUST be a list of strings, containing the names of the relations you have selected. Example Input: Question: "who was the president after jfk died" Candidate Relations: {"government.president", "government.president.successor", "location.location.containedby", "people.person.place_of_birth"} K: 2 Output:
@@
Appendix F The Use of Large Language Models (LLMs)
Large language models (LLMs) were only used to improve the clarity, grammar, and fluency of the manuscript. They were not involved in the development of research ideas, experimental design, data analysis, or any other aspect of the scientific content.